# Super approximation for $\text{SL}_{2}\times\text{SL}_{2}$ and $\text{ASL}_{2}$

### v3

@@ 5.1. A large set of commuting elements
$W(n):=\{\gamma\in\Gamma:\text{tr}(\gamma)=n(\operatorname{mod}Q_{1})\}.$
By Proposition 4.1, for any $n\in\mathbb{Z}$,
$\nu^{2l}(W(n))<Q_{1}^{-c_{1}}.$ (5.9)
If we take
${\delta}<{\frac{\rho_{2}c_{1}}{12}},$ (5.10)
then $\nu^{(2l)}(A_{3})>Q_{1}^{-\frac{c_{1}}{2}}$, and together with (5.9), it implies that there exists a set $A_{4}\subset A_{3}$, such that
$|A_{4}(\operatorname{mod}Q_{1})|>Q_{1}^{c_{1}/2}$ (5.11)
$\forall\gamma_{1},\gamma_{2}\in A_{4},\gamma_{1}\neq\gamma_{2}\Rightarrow\text{tr}(\gamma_{1})\neq\text{tr}(\gamma_{2})(\text{mod }Q_{1})$ (5.12)
Next we need to use our assumption (5.2) to get a control over the growth of $\pi_{Q_{1}}(A_{1})$. By Lemma 2.2 from [Hel08], (5.2) implies for any $l\geq 3$,
$|\pi_{q}(\prod_{l}A)|\leq\left(\frac{|\pi_{q}(A\cdot A\cdot A)|}{|\pi_{q}(A)|}\right)^{l-2}|\pi_{q}(A)|$ (5.13)
$<$ $|\pi_{q}(A)|^{1+\delta(l-2)}.$ (5.14)
It follows from Proposition 4.1 that for any $\gamma_{0}\in\Lambda_{q}$,
$\nu^{(2l)}\{\gamma\in\Lambda:\gamma\equiv\gamma_{0}(\operatorname{mod}Q_{1})\}<Q_{1}^{-c_{1}}.$ (5.15)
From (5.15) and (5.10), we conclude that
$|\pi_{Q_{1}}(A_{1})|>Q_{1}^{c_{1}}q^{-3\delta}>Q_{1}^{\frac{c_{1}}{2}}.$
Next we show small expansion for $A$ implies small expansion for $A_{1}$:
$|\pi_{Q_{1}}(A_{1}\cdot A_{1}\cdot A_{1})|<|\pi_{Q_{1}}(A_{1})|^{1+\frac{4\delta}{\rho_{2}c_{1}}}.$ (5.16)
To see this, suppose not, then
$|\pi_{q}(A^{4})|\geq$ $|\pi_{Q_{1}}(A_{1}\cdot A_{1}\cdot A_{1})|\cdot\max_{\gamma_{0}\in\Lambda_{Q_{1}}}|\pi_{Q_{1}}(A)\cap\{(\gamma_{0},\gamma):\gamma\in\text{SL}_{2}(\mathbb{Z}/q\mathbb{Z})\}|$
$\geq$ $|\pi_{Q_{1}}(A_{1})|^{1+\frac{4\delta}{\rho_{2}c_{1}}}\cdot\max_{\gamma_{0}\in\Lambda_{Q_{1}}}|\pi_{Q_{1}}(A)\cap\{(\gamma_{0},\gamma):\gamma\in\text{SL}_{2}(\mathbb{Z}/q\mathbb{Z})\}|$
$\geq$ $|\pi_{Q_{1}}(A_{1})|^{\frac{4\delta}{\rho_{2}c_{1}}}\cdot|\pi_{q}(A)|$
$>$ $|\pi_{q}(A)|^{1+\delta}$ (5.17)
which contradicts (5.2).
For each $\alpha\in A_{4}$, let $C_{\alpha}=\{\gamma\alpha\gamma^{-1}:\gamma\in A_{1}\}$. Since distinct elements in $A_{4}$ have distinct eigenvalues mod $Q_{1}$, the sets $\{\pi_{Q_{1}}(C_{\alpha})\}_{\alpha\in A_{1}}$ are mutually disjoint. Therefore,
$\sum_{\alpha\in A_{4}}|\pi_{Q_{1}}(C_{\alpha})|=|\cup_{\alpha\in A_{4}}\pi_{Q_{1}}(C_{\alpha})|<|\pi_{Q_{1}}(A_{1}^{3})|<|\pi_{Q_{1}}(A_{1})|^{1+\frac{4\delta}{\rho_{2}c_{1}}}.$
Thus, for some $\alpha_{0}=\left(\begin{matrix}\alpha_{11}&\alpha_{12}\\ \alpha_{21}&\alpha_{22}\end{matrix}\right)\in A_{4}$,
$|\pi_{Q_{1}}(C_{\alpha_{0}})|\leq|\pi_{Q_{1}}(A)|^{1+\frac{4\delta}{\rho_{2}c_{1}}}Q_{1}^{-c_{1}/2}<|\pi_{Q_{1}}(A)|Q_{1}^{-\frac{c_{1}}{4}}.$ (5.18)
provided that
$\delta<\frac{\rho_{2}c_{1}^{2}}{48}.$ (5.19)
To proceed, we first diagonalise $\alpha_{0}$ in an extended number field, i.e., let
be the two eigenvalues of $\alpha_{0}$. Let $K=\mathbb{Q}[\lambda_{1}]$, and so $\mathcal{O}=\mathcal{O}_{K}=\mathbb{Z}[\lambda_{1}]$. Let $M=\left(\begin{matrix}\lambda_{1}-\alpha_{22}&\lambda_{2}-\alpha_{22}\\ \alpha_{21}&\alpha_{21}\end{matrix}\right)$. Then
$M^{-1}\alpha_{0}M=\left(\begin{matrix}\lambda_{1}&0\\ 0&\lambda_{2}\end{matrix}\right).$
Since $\alpha_{0}$ satisfies the condition (5.6) for $\gamma$,
$\lambda_{1}\not\equiv\lambda_{2}(\operatorname{mod}p^{[4\rho_{1}n]})$ (5.20)
for any $p^{n}||q_{1}$.
From (5.18), there exists $x_{0}\in A_{4}$ and
$A_{5}:=\{x\in A:x\alpha_{0}x^{-1}=x_{0}\alpha_{0}x_{0}^{-1}(\operatorname{mod}Q_{1})\},$ (5.21)
such that
$|\pi_{Q_{1}}(A_{5})|>Q_{1}^{\frac{c_{1}}{4}}.$ (5.22)
Therefore, $A_{5}^{-1}\cdot x_{0}$ commute with $\alpha_{0}(\operatorname{mod}Q_{1})$ and
$|\pi_{Q_{1}}(A_{5}^{-1}\cdot x_{0})|>Q_{1}^{\frac{c_{1}}{4}}.$ (5.23)
Let $C_{1},C_{2},C_{3}$ be the implied constants of Theorem B for $\delta=\frac{c_{1}}{17}$.
The condition (5.23) implies that there exists $\upsilon\in[\frac{c_{1}}{24},\frac{1}{2}]$, such that
$|\pi_{Q_{1}^{\{\upsilon(1+\frac{1}{C_{1}})\}}}(A_{5})|>|\pi_{Q_{1}^{\{\upsilon\}}}(A_{5})|Q_{1}^{\frac{\upsilon c_{1}}{8C_{1}}}.$ (5.24)
If we let $A_{6}=\left((A_{5}^{-1}\cdot x_{0})\cdot(A_{5}^{-1}\cdot x_{0})^{-1}\right)\cap\Gamma(Q_{1}^{\{\upsilon\}})=A_{5}^{-1}A_{5}\cap\Gamma(Q_{1}^{\{\upsilon\}})$, then
$|\pi_{Q_{1}^{\{\upsilon(1+\frac{1}{C_{1}})\}}}(A_{6})|>Q_{1}^{\frac{\upsilon c_{1}}{8C_{1}}}.$ (5.25)
Let
$Q_{2}:=q_{1}^{\{\rho_{2}\upsilon-9\rho_{1}\}},$
where we will take $\rho_{2}\upsilon\gg 9\rho_{1}$. Since $\text{Det}(M)=\alpha_{21}\sqrt{\text{tr}(\alpha_{0})^{2}-4}$, from (5.7), (5.6), we have
$p^{[6\rho_{1}n]}\nmid\text{Det}(M),\forall p^{n}||q_{1}.$ (5.26)
Combining (5.26) with (5.20), we conclude that the matrix $M$ conjugate every matrix $\sigma$ in $A_{6}$ to a matrix in $\text{SL}_{2}(\mathcal{O})$ which is congruent to identity mod $Q_{2}$ and diagonal mod $Q_{2}^{\left\{1+\frac{1}{C_{1}}\right\}}$. In other words, there exists $a_{\sigma}\in\mathcal{O}_{K}$ and $a_{\sigma}\equiv 1(\operatorname{mod}Q_{2})$ such that
$M^{-1}\sigma M\equiv\left(\begin{matrix}a_{\sigma}&0\\ 0&a_{\sigma}^{-1}\end{matrix}\right)(\operatorname{mod}Q_{2}^{\left\{1+\frac{1}{C_{1}}\right\}}).$ (5.27)
It also follows from (5.25) that
$\Big{|}\pi_{Q_{2}^{\left\{1+\frac{1}{C_{1}}\right\}}}(A_{6})\Big{|}>Q_{2}^{\frac{c_{1}}{16C_{1}}}.$ (5.28)
if
$\rho_{1}<\frac{\upsilon c_{1}}{2^{5}3^{3}}\rho_{2},$ (5.29)
which is guaranteed if
$\rho_{1}<\frac{c_{1}^{2}}{2^{8}3^{4}}\rho_{2}.$ (5.30)
For each $\gamma\in A_{6}$, let ${q}_{\gamma}$ be the largest divisor of $Q_{2}^{\left\{1+\frac{1}{C_{1}}\right\}}$ such that $\gamma\equiv 1(\operatorname{mod}q_{\gamma})$. Obviously, $Q_{2}|q_{\gamma}$. Then we can write
$A_{6}=\sqcup_{Q_{2}|q^{\prime}|Q_{2}^{\left\{1+\frac{1}{C_{1}}\right\}}}B_{q^{\prime}},$
where $B_{q^{\prime}}=\{\gamma\in A_{6}:q_{\gamma}=q^{\prime}\}$. Since the number of divisors of $Q_{2}^{\left\{1+\frac{1}{C_{1}}\right\}}$ is bounded by $q^{0+}$, there exists $\tilde{q}$ such that $Q_{2}|\tilde{q}|Q_{2}^{\left\{1+\frac{1}{C_{1}}\right\}}$ and
$\Big{|}B_{\tilde{q}}(\operatorname{mod}Q_{2}^{\left\{1+\frac{1}{C_{1}}\right\}})\Big{|}>Q_{2}^{\frac{c_{1}}{16C_{1}}-}.$ (5.31)
5.2. Applying sum-product
Let $H=\{h_{\sigma}=(a_{\sigma}-a_{\sigma}^{-1})/\tilde{q}:\sigma\in B_{\tilde{q}}\}\subset\mathcal{O}$, where $a_{\sigma}$ is as given in (5.27). Since the map $a_{\sigma}\mapsto a_{\sigma}-a_{\sigma}^{-1}$ is bounded to one, we have
$\Big{|}H(\operatorname{mod}{Q_{2}^{\{\frac{1}{C_{1}}\}}})\Big{|}>Q_{2}^{\frac{c_{1}}{16C_{1}}-}.$ (5.32)
Recall $\mathcal{O}$ is a degree 2 extension of $\mathbb{Z}$. We apply Theorem B to the set $H$ with $\mathfrak{a}=(Q_{2}^{\{\frac{1}{C_{1}}\}})$ and $\delta=\frac{c_{1}}{17}$. Theorem B then produces some $\xi\in\mathcal{O}$, $\mathfrak{a}^{\prime}\supset\mathfrak{a}^{C_{1}}$, $c_{3}=c_{3}(c_{1})>0$ such that
$\sum_{C_{3}}H^{C_{2}}\supset\mathbb{Z}\xi(\operatorname{mod}\mathfrak{a}^{\prime}),$ (5.33)
and
$|\mathbb{Z}\xi(\operatorname{mod}\mathfrak{a}^{\prime})|>N(\mathfrak{a})^{c_{3}}>q_{1}^{\frac{\rho_{2}\upsilon c_{3}}{C_{1}}}>q_{1}^{\frac{\rho_{2}c_{1}c_{3}}{24C_{1}}}.$ (5.34)
We pass (5.34) to a local property. From (5.34), it follows that there exists $q_{2}||q_{1}$, $q_{2}>q_{1}^{\frac{c_{3}}{2C_{1}}}$, such that for any $p^{n}||q_{2}$, letting $\overline{\mathfrak{a}_{p}}=\langle\mathfrak{a}^{\prime},p^{n}\rangle$, the ideal of $\mathcal{O}$ generated by $\mathfrak{a}^{\prime}$ and $p^{n}$, we have
$|\mathbb{Z}\xi(\operatorname{mod}\overline{\mathfrak{a}_{p}})|>p^{\frac{\rho_{2}c_{1}c_{3}n}{48C_{1}}}.$
Since there are at most two prime ideals in $\mathcal{O}$ lying above $p$, there exists a prime ideal $\mathcal{P}_{p}\supset(p)$ in $\mathcal{O}$, $\mathcal{P}_{p}^{m_{p}}||\bar{\mathfrak{a}_{p}}$, such that
$|\mathbb{Z}\xi(\operatorname{mod}\mathcal{P}_{p}^{m_{p}})|>p^{\frac{\rho_{2}c_{1}c_{3}n}{96C_{1}}}.$ (5.35)
For simplicity, we assume for each $p|q_{3}$, the arising prime ideal $\mathcal{P}_{p}$ is unramified, so $v_{p}(\tilde{q})=v_{\mathcal{P}}(\tilde{q})$. The general case adds no additional technical challenge.
We obtain from (5.35), $\mathfrak{a}^{\prime}\supset\mathfrak{a}^{C_{1}}$ and the definition of $\mathfrak{a}$ that
$\frac{\rho_{2}c_{1}c_{3}}{96C_{1}}n<m_{p}<\rho_{2}n.$ (5.36)
It also follows from (5.35) that for any $p|q_{2}$,
$v_{p}(Q_{2})\leq v_{p}(\tilde{q})<v_{p}(Q_{2}^{\{1+\frac{1}{C_{1}}\}}).$ (5.37)
5.3. Constructing an appropriate element $\zeta$
We also need to find an element $\zeta=\left(\begin{matrix}a&b\\ c&d\end{matrix}\right)$ from $A_{1}$ which is very non-commutative with $\alpha_{0}=\left(\begin{matrix}\alpha_{11}&\alpha_{12}\\ \alpha_{21}&\alpha_{22}\end{matrix}\right)$. This leads us to work with the linear form
$\mathcal{L}_{1}\left(\left(\begin{matrix}a&b\\ c&d\end{matrix}\right)\right)=\alpha_{21}a+(\alpha_{22}-\alpha_{11})c-\alpha_{21}d,$
which gives the 2-1 entry of $\alpha_{0}\zeta-\zeta\alpha_{0}$. If $\mathcal{L}_{1}(\zeta)$ is non-zero, this certainly implies $\zeta$ and $\alpha_{0}$ are not commutative.
Let
$\mathcal{E}_{3}(Q)=\{\gamma\in\text{SL}_{2}(\mathbb{Z}):\mathcal{L}_{1}(\gamma)\equiv 0(\operatorname{mod}Q)\}.$
By Proposition 4.1, for any $Q|q_{2}$ ,
@@ 5.3. Constructing an appropriate element $\zeta$
$A_{7}=A_{1}-\cup_{\begin{subarray}{c}Q|q_{2}\\ Q>q_{2}^{\rho_{1}}\end{subarray}}\mathcal{E}_{3}(Q).$
By (5.4), $\nu^{(2l)}(A_{1})>q^{-3\delta}$. We also have $q_{2}>q_{1}^{\frac{c_{3}}{2C_{1}}}>q^{\frac{c_{3}}{4C_{1}}}$. Then
$\nu^{(2l)}(A_{7})>q^{-4\delta}$
if $q_{2}^{-\rho_{1}c_{1}}<q^{-4\delta}$, or
${\delta}<\frac{c_{1}c_{3}\rho_{1}}{16C_{1}}.$ (5.39)
Then we produce some $q_{3}||q_{2}$, $q_{3}>q_{2}^{\frac{1}{2}}$, $\zeta_{1}\in A_{1}$ such that for any $p^{n}||q_{3}$,
$L({\zeta_{1}})\neq 0(\operatorname{mod}p^{[2\rho_{1}n]}).$ (5.40)
From (5.40) and (5.20), we have $M\zeta_{1}M^{-1}=\left(\begin{matrix}a_{11}&a_{12}\\ a_{21}&a_{22}\end{matrix}\right)$ with $a_{12}$ or $a_{21}\not\equiv 0(\operatorname{mod}p^{[6n\rho_{1}]})$. Take any element $\zeta_{2}\in B_{\tilde{q}}$. The second inequality of (5.37) ensures that $v_{p}(\zeta_{2})=v_{p}(\tilde{q})$.
If we let $\zeta=\zeta_{1}\zeta_{2}\zeta_{1}^{-1}$ and write $M{\zeta}M^{-1}=\left(\begin{matrix}\zeta_{11},\zeta_{12}\\ \zeta_{21},\zeta_{22}\end{matrix}\right)$. Then
$v_{p}(M\zeta M^{-1}-1)\geq v_{p}(\tilde{q})-[6\rho_{1}n]$ (5.41)
@@
$\min\{v_{p}(\zeta_{12}),v_{p}(\zeta_{21})\}\leq v_{p}(\tilde{q})+[10\rho_{1}n].$ (5.42)
5.4. Creating a line
Let
$\left\{\vec{e}=\left(\begin{matrix}0&1\\ 0&0\end{matrix}\right),\vec{f}=\left(\begin{matrix}0&0\\ 1&0\end{matrix}\right),\vec{h}=\left(\begin{matrix}1&0\\ 0&-1\end{matrix}\right)\right\}$
[…]
To proceed, we recall the following well known identity.
@@ 5.4. Creating a line
Lemma 5.43 has an obvious generalization to a general ideal by multiplicativity.
To apply Lemma 5.43, we take $x=M^{-1}\sigma M$ given at (5.27) and $y=M^{-1}\zeta M$, with corresponding moduli $\mathfrak{q}_{1}=\prod_{p|q_{3}}\mathcal{P}_{p}^{v_{\mathcal{P}}(\tilde{q})}$ and $\mathfrak{q}_{2}=\prod_{p|q_{3}}\mathcal{P}_{p}^{v_{\mathcal{P}}(\tilde{q})-[6\rho_{1}n]}$.
Then
$M^{-1}\sigma\zeta\sigma^{-1}\zeta^{-1}M\equiv 1(\operatorname{mod}\mathfrak{q}_{1}\mathfrak{q}_{2})$ (5.44)
@@
$M^{-1}\sigma\zeta\sigma^{-1}\zeta^{-1}M\equiv 1+\tilde{q}h_{\sigma}\left(\zeta_{21}\vec{e}-\zeta_{12}\vec{f}\right)(\operatorname{mod}\mathfrak{q}_{1}\mathfrak{q}_{2}^{2})$ (5.45)
Applying Lemma 5.43 iteratively, for any $\sigma_{1},\sigma_{2},\cdots,\sigma_{C_{3}}\in A_{6}$, we have
$M^{-1}[\sigma_{C_{3}},\cdots[\sigma_{2},[\sigma_{1},\zeta]]\cdots]M\equiv 1(\operatorname{mod}(\mathfrak{q}_{1}^{C_{3}-1}\mathfrak{q}_{2}^{2})$ (5.46)

### v4

@@ 5.4. Constructing a line
$\mathcal{E}_{3}(Q)=\{\gamma\in\text{SL}_{2}(\mathbb{Z}):\mathcal{L}(\gamma)\equiv 0(\operatorname{mod}Q)\}.$
By Proposition 4.1, for any sufficiently large $Q|q_{2}$ ,
@@
$A_{3}=A_{0}-\cup_{\begin{subarray}{c}Q|q_{2}\\ Q>q_{2}^{\rho_{1}}\end{subarray}}\mathcal{E}_{3}(Q).$
By (5.1), (5.49) and (5.57),
$\nu^{(2l)}(A_{3})>q^{-3\delta}-q_{2}^{-\rho_{1}c_{1}-}>q^{-3\delta}-q_{1}^{-\frac{\varepsilon_{0}\rho_{1}c_{1}}{4}-}>q^{-3\delta}-q^{-\frac{\varepsilon_{0}\rho_{1}c_{1}}{8}-}>q^{-4\delta},$
if taking $L_{1}$ sufficiently large.
Then using the same argument for constructing $A_{1}$ in Lemma 5.7, by taking $L_{1}$ sufficiently large, we produce some $q_{3}\|q_{2}$, $q_{3}>q_{2}^{\frac{1}{2}}$, $\zeta_{1}\in A_{3}$ such that for any $p^{n}\|q_{3}$,
$\mathcal{L}({\zeta_{1}})\neq 0(\operatorname{mod}p^{[2\rho_{1}n]}),$
which of course implies for any $p^{n}\|q_{3}$,
$\mathcal{L}({\zeta_{1}})\neq 0(\operatorname{mod}\mathcal{P}_{p}^{\{O(\rho_{1}n)\}}).$ (5.58)
The matrix $\zeta_{1}$ itself is very non-commutative with $\alpha_{0}$ in the sense of (5.58). However for our purpose, we need to find a non-commutative element that is also sufficiently close to identity (at the scale $w_{1}$) as required by lemma 5.53, and then conjugate it by $M$.
The modification is as follows: First recall the diagonalization of $\alpha_{0}$ at (5.30) and that $\lambda_{1}-\lambda_{2}=\sqrt{\text{tr}(\alpha_{0})^{2}-4}$. For any $p^{n}\|q_{3}$, since $v_{\mathcal{P}_{p}}(\text{tr}(\alpha_{0})^{2}-4)=O(\rho_{1}n)$, we have
$\lambda_{1}\not\equiv\lambda_{2}(\operatorname{mod}\mathcal{P}_{p}^{\{O(\rho_{1}n)\}}).$ (5.59)
From (5.58) and (5.59), we have
$\zeta_{2}=M^{-1}\zeta_{1}M=\left(\begin{matrix}a_{11}&a_{12}\\ a_{21}&a_{22}\end{matrix}\right)$
with $a_{12}$ or $a_{21}\not\equiv 0(\operatorname{mod}\mathcal{P}_{p}^{\{O(\rho_{1}n)\}})$. Therefore, there exists an exact divisor $q_{4}\|q_{3},q_{4}\geq q_{3}^{\frac{1}{2}}$ such that one of $a_{12}$ and $a_{21}$, let us say $a_{21}$ without loss of generality, satisfies: $\forall p^{n}\|q_{4}$,
$a_{21}\not\equiv 0(\operatorname{mod}\mathcal{P}_{p}^{\{O(n\rho_{1})\}}).$ (5.60)
Take any element $\bar{h}\in H$ as given by Lemma 5.32 with $\bar{h}\equiv\left(\begin{matrix}a_{\bar{h}}&0\\ 0&(a_{\bar{h}})^{-1}\end{matrix}\right)(\operatorname{mod}q_{4}^{\{w_{1}+w_{2}\}})$. A direct computation gives
$\zeta=[\zeta_{2},\bar{h}]=\zeta_{2}\bar{h}\zeta_{2}^{-1}{\bar{h}}^{-1}\equiv 1+\left(\begin{matrix}a_{12}a_{21}(1-(a_{\bar{h}})^{-2})&a_{11}a_{12}(1-(a_{\bar{h}})^{2})\\ a_{21}a_{22}(1-(a_{\bar{h}})^{-2})&a_{12}a_{21}(1-(a_{\bar{h}})^{2})\end{matrix}\right)(\operatorname{mod}q_{4}^{\{w_{1}+w_{2}-O(\rho_{1})\}}),$ (5.61)
where again we have a loss of an insignificant $O(\rho_{1})$ term in the exponent of $q_{4}$ because of the denominator of $M$.
We have the following estimate for $v_{\mathcal{P}_{p}}(\bar{h}-1)$:
$v_{\mathcal{P}_{p}}(q_{4}^{\{w_{1}\}})\leq v_{\mathcal{P}_{p}}(\bar{h}-1)<v_{\mathcal{P}_{p}}(q_{4}^{\{w_{1}+\frac{w_{2}}{C_{1}}\}}).$ (5.62)
The second inequality of (5.62) in particular holds, because otherwise, for any $h\in H$,
$v_{\mathcal{P}_{p}}(\lambda_{h})=v_{\mathcal{P}_{p}}(\lambda_{\bar{h}})\geq v_{\mathcal{P}_{p}}(q_{4}^{\{\frac{w_{2}}{C_{1}}\}})=v_{\mathcal{P}_{p}}(q_{1}^{\{\frac{w_{2}}{C_{1}}\}}),$ (5.63)
recalling (5.43) and Property (iv) in Lemma 5.32. Following the proof of Theorem B in [TZ23], (5.63) would imply the implied arithmetic progression $\mathbb{Z}\xi$ has no contribution at the modulus $\mathcal{Q}_{p}$, i.e., $\mathbb{Z}\xi(\operatorname{mod}\mathcal{Q}_{p})=\{0\}$. This contradicts the thickness bound of the projection of the arithmetic progression $\mathbb{Z}\xi$ at $\mathcal{Q}_{p}$ in the definition of $q_{2}$ at (5.48). So (5.62) holds.
Let $\mathscr{Q}_{1}$ be the ideal of $\mathcal{O}$ generated by $q_{4}$ and all entries of $\xi-1$.
Write
$\zeta\equiv 1+\mathfrak{q}_{1}\left(\begin{matrix}X&Y\\ Z&-X\end{matrix}\right)(\operatorname{mod}\mathscr{Q}_{1}^{2}),$
where $\mathfrak{q}_{1}$ is an element in $\mathcal{O}$ that uniformizes $\mathscr{Q}_{1}$, and $\left(\begin{matrix}X&Y\\ Z&-X\end{matrix}\right)$ is a primitive matrix in $\text{Mat}_{2}(\mathcal{O})$. From (5.61) and (5.60), we have for any $p^{n}\|q_{4}$,
$v_{\mathcal{P}_{p}}(\mathscr{Q}_{1})=v_{\mathcal{P}_{p}}(\mathfrak{q}_{1})=v_{\mathcal{P}_{p}}(\bar{h}-1)\pm O(\rho_{1}n)$ (5.64)
@@
$v_{\mathcal{P}_{p}}(Z)=O(\rho_{1}n).$ (5.65)
Requiring $C_{1}>100$ and taking $L_{2}$ sufficiently large then give (5.54) in Lemma 5.53.
To proceed, we recall the following well known identity (see Lemma 6.1, [BG09]).
@@
Lemma 5.66 has an obvious generalization to general ideals by multiplicativity:
Lemma 5.67.
Let $\mathcal{I}_{1},\mathcal{I}_{2}$ be two ideals of $\mathcal{O}$, and let $x,y\in\text{SL}_{2}(\mathcal{O})$, $x\equiv 1(\operatorname{mod}\mathcal{I}_{1}),y\equiv 1(\operatorname{mod}\mathcal{I}_{2})$. Then
$xyx^{-1}y^{-1}\equiv 1(\operatorname{mod}\mathcal{I}_{1}\mathcal{I}_{2}),$
@@
$xyx^{-1}y^{-1}\equiv 1+xy-yx(\operatorname{mod}\mathcal{I}_{1}\mathcal{I}_{2}\langle\mathcal{I}_{1},\mathcal{I}_{2}\rangle).$
To apply Lemma 5.67, we take $x=\zeta$ given by Lemma 5.53 with $\mathcal{I}_{1}=\mathscr{Q}_{1}$, and $y=h_{1}\in H$ given in Lemma 5.32 with $\mathcal{I}_{2}=\mathscr{Q}_{2}:=(q_{4}^{\{w_{1}\}})$.
Write $\mathscr{Q}_{3}=(q_{4}^{\{w_{2}-O(\rho_{1})\}})$, where the implied constant for $O(\rho_{1})$ is taken so that $\mathscr{Q}_{3}$ divides $\langle\mathscr{Q}_{1},\mathscr{Q}_{2}\rangle$, noticing that for each $p^{n}\|q_{4}$, $v_{\mathcal{P}_{p}}(\langle\mathscr{Q}_{1},\mathscr{Q}_{2}\rangle)=v_{\mathcal{P}_{p}}(q_{4}^{\{w_{1}\}})\pm O(\rho_{1}n)$.
Recall $h_{1}\equiv\left(\begin{matrix}a_{h_{1}}&0\\ 0&a_{h_{1}}^{-1}\end{matrix}\right)(\operatorname{mod}\mathscr{Q}_{2}\mathscr{Q}_{3})$ and $a_{h_{1}}-a_{h_{1}}^{-1}=q_{1}^{\{w_{1}\}}\lambda_{h_{1}}$. Also write $\mathfrak{q}_{2}=q_{1}^{\{w_{1}\}}$, which is a uniformizer for $(q_{4}^{\{w_{1}\}})$. Then from Lemma 5.67,
$\zeta^{\prime}:=\zeta h_{1}\zeta^{-1}h_{1}^{-1}\equiv 1(\operatorname{mod}\mathscr{Q}_{1}\mathscr{Q}_{2})$ (5.68)